The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34in., find the dimensions of the rectangle.
Answer by uma(370)

Let the width be x cm
Then length measures = 2x+2
The perimeter given to be 34
==> 2[x+2x+2) = 34
==> 2[3x+2] = 34
==> 6x + 4 = 34
==> 6x = 34 - 4
==> 6x = 30
==> x = 5
Thus the width = 5 in
and length = 2(5)+2
= 10+2
= 12 in

Good Luck!!!!